# Solving Football by Exploiting Equilibrium Structure of 2p0s Differential Games with One-Sided Information

### v3

@@ Appendix F Analytical Examples
In this section, we walk through the derivation of analytical NEs for two problems: Hexner’s game and a zero-sum variant of the classic beer-quiche game. The former is differential where players take actions simultaneously; the latter is dynamic and turn-based. Both have one-sided payoff information and finite time horizons. These examples are reproduced from Ghimire et al. (2024) with permission.
@@ F.1 Hexner’s Game: Analytical Solution
Here we discuss the solution to Hexner’s game using primal and dual formulations (i.e., $\text{P}_{1}$ and $\text{P}_{2}$) on a differential game as proposed in Hexner (1979). Consider two players with linear dynamics
@@
$\qquad\Bigl.\|x_{1}(T)-z\theta\|^{2}_{K_{1}(T)}-\|x_{2}(T)-z\theta\|^{2}_{K_{2}(T)}\Bigr],$
where, $z\in\mathbb{R}^{d_{x}}$. $R_{1}$ and $R_{2}$ are continuous, positive-definite matrix-valued functions, and $K_{1}(T)$ and $K_{2}(T)$ are positive semi-definite matrices. All parameters are publicly known except for $\theta$, which remains private. Player 1’s objective is to get closer to the target $z\theta$ than Player 2. However, since Player 2 can deduce $\theta$ indirectly through Player 1’s control actions, Player 1 may initially employ a non-revealing strategy. This involves acting as though he only knows about the prior distribution $p_{\theta}$ (rather than the true $\theta$) to hide the information, before eventually revealing $\theta$.
@@
The following properties are necessary (see Ghimire et al. (2024) for details):
@@ F.2 Example of a Turn-Based Game
We present a zero-sum variant of the classic beer-quiche game, which is a turn-based incomplete-information game with a perfect Bayesian equilibrium. Unlike in Hexner’s game, Player 1 in beer-quiche game wants to maximize his payoff, and Player 2 wants to minimize it; hence, Vex becomes a Cav.
@@ G.3 OpenSpiel Implementations and Hyperparameters
We use OpenSpiel (Lanctot et al., 2019), a collection of various environments and algorithms for solving single and multi-agent games. We select OpenSpiel due to its ease of access and availability of wide range of algorithms. The first step is to write the game environment with simultaneous moves for the stage-game and the multi-stage games (with 4 decision nodes). Note that to learn the policy, the algorithms in OpenSpiel require conversion from simultaneous to sequential game, which can be done with a built-in method.
In the single-stage game, P1 has two information states representing his type, and P2 has only one information state (i.e., the starting position of the game which is fixed). In the case of the 4-stage game, the information state (or infostate) is a vector consisting of the P1’s type (2-D: [0, 1] for type-1, [1, 0] for type-2), states of the players (8-D) and actions of the players at each time step $(4\times 2\times U)$. The 2-D “type” vector for P2 is populated with 0 as she has no access to P1’s type. For example, the infostate at the final decision node for a type-1 P1 could be $0,1,x^{(8)},\mathbbm{1}_{u_{0}}^{(U)},\mathbbm{1}_{d_{0}}^{(U)},\cdots,\mathbbm{1}_{d_{2}}^{(U)},\bm{0}^{(U)},\bm{0}^{(U)}]$, and $0,0,x^{(8)},\mathbbm{1}_{u_{0}}^{(U)},\mathbbm{1}_{d_{0}}^{(U)},\cdots,\mathbbm{1}_{d_{2}}^{(U)},\bm{0}^{(U)},\bm{0}^{(U)}]$ for P2, where $u_{k}$, $d_{k}$ represent the index of the actions at $k^{th}$ decision node, $k=0,1,2,3$
@@ G.4 Joint-Perturbation Simultaneous Pseudo-Gradient (JPSPG)
The core idea in the JPSPG algorithm is the use of pseudo-gradient instead of computing the actual gradient of the utility to update players’ strategies. By perturbing the parameters of a utility function (which consists of the strategy), an unbiased estimator of the gradient of a smoothed version of the original utility function is obtained. Computing pseudo-gradient can often be cheaper as faster than computing exact gradient, and at the same time suitable in scenarios where the utility (or objective) functions are "black-box" or unknown. Building on top of pseudo-gradient, Martin & Sandholm (2024) proposed a method that estimates the pseudo-gradient with respect to all players’ strategies simultaneously. The implication of this is that instead of multiple calls to estimate the pseudo-gradient, we can estimate the pseudo-gradient in a single evaluation. More formally, let $\textbf{f}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ be a vector-valued function. Then, its smoothed version is defined as:
@@
In games with discrete-action spaces, where strategy is the probability distribution over the actions, JPSPG can be directly applied to get mixed strategy. However, for continuous-action games, a standard implementation would result in pure strategy solution than mixed. In order to compute a mixed strategy, we can turn into neural network as a strategy with an added randomness that can be learned as described in Martin & Sandholm (2023, 2024). We similarly define two strategy networks for each player, the outputs of which are scaled based on the respective action bounds with the help of hyperbolic tangent (tanh) activation on the final layer. The input to the strategy networks (a single hidden layered neural network with 64 neurons and output neuron of action-space dimension) are the state of the player and a random variable whose mean and variance are trainable parameters. We follow the architecture as outlined by Martin & Sandholm (2024) in their implementation of continuous-action Goofspiel. We would like to thank the authors for providing an example implementation of JPSPG on a normal-form game.
@@ G.6 Value Network Training Details
We briefly discuss the training procedure of the value networks. As mentioned in the main paper, both the primal and the dual value functions are convex with respect to $p$ and $\hat{p}$ respectively. As a result, we use Input Convex Neural Networks (ICNN) (Amos et al., 2017) as the neural network architecture. Starting from $T-\tau$, solutions of the optimization problem $\text{P}_{1}$ for sampled $(X,p)$ is saved and the convex value network is fit to the saved training data. The model parameters are saved and are then used in the optimization step at $T-2\tau$. This is repeated until the value function at $t=0$ is fit. The inputs to the primal value network are the joint states containing position and velocities of the players $X$ and the belief $p$.
@@
Appendix H Solving Hexner’s game via MPC
@@
Appendix I A Differentiable 11-vs-11 American Football Game

### v4

@@ Appendix F Analytical Examples
In this section, we walk through the derivation of analytical NEs for two problems: Hexner’s game and a zero-sum variant of the classic beer-quiche game. The former is differential where players take actions simultaneously; the latter is dynamic and turn-based. Both have one-sided payoff information and finite time horizons. These examples are reproduced from ghimire24a with permission.
@@ F.1 Hexner’s Game: Analytical Solution
Here we discuss the solution to Hexner’s game using primal and dual formulations (i.e., $\text{P}_{1}$ and $\text{P}_{2}$) on a differential game as proposed in hexner1979differential. Consider two players with linear dynamics
@@
$\quad\Bigl.\|x_{1}(T)-z\theta\|^{2}_{K_{1}(T)}-\|x_{2}(T)-z\theta\|^{2}_{K_{2}(T)}\Bigr],$
where, $z\in\mathbb{R}^{d_{x}}$. $R_{1}$ and $R_{2}$ are continuous, positive-definite matrix-valued functions, and $K_{1}(T)$ and $K_{2}(T)$ are positive semi-definite matrices. All parameters are publicly known except for $\theta$, which remains private. Player 1’s objective is to get closer to the target $z\theta$ than Player 2. However, since Player 2 can deduce $\theta$ indirectly through Player 1’s control actions, Player 1 may initially employ a non-revealing strategy. This involves acting as though it only knows about the prior distribution $p_{\theta}$ (rather than the true $\theta$) to hide the information, before eventually revealing $\theta$.
@@
The following properties are necessary (see ghimire24a for details):
@@ F.2 Example of a Turn-Based Game
We present a zero-sum variant of the classic beer-quiche game, which is a turn-based incomplete-information game with a perfect Bayesian equilibrium. Unlike in Hexner’s game, Player 1 in beer-quiche game wants to maximize its payoff, and Player 2 wants to minimize it; hence, Vex becomes a Cav.
@@ G.3 OpenSpiel Implementations and Hyperparameters
We use OpenSpiel (LanctotEtAl2019OpenSpiel), a collection of various environments and algorithms for solving single and multi-agent games. We select OpenSpiel due to its ease of access and availability of wide range of algorithms. The first step is to write the game environment with simultaneous moves for the stage-game and the multi-stage games (with 4 decision nodes). Note that to learn the policy, the algorithms in OpenSpiel require conversion from simultaneous to sequential game, which can be done with a built-in method.
In the single-stage game, P1 has two information states representing its type, and P2 has only one information state (i.e., the starting position of the game which is fixed). In the case of the 4-stage game, the information state (or infostate) is a vector consisting of the P1’s type (2-D: [0, 1] for type-1, [1, 0] for type-2), states of the players (8-D) and actions of the players at each time step $(4\times 2\times U)$. The 2-D “type” vector for P2 is populated with 0 as it has no access to P1’s type. For example, the infostate at the final decision node for a type-1 P1 could be $0,1,x^{(8)},\mathbbm{1}_{u_{0}}^{(U)},\mathbbm{1}_{d_{0}}^{(U)},\cdots,\mathbbm{1}_{d_{2}}^{(U)},\bm{0}^{(U)},\bm{0}^{(U)}]$, and $0,0,x^{(8)},\mathbbm{1}_{u_{0}}^{(U)},\mathbbm{1}_{d_{0}}^{(U)},\cdots,\mathbbm{1}_{d_{2}}^{(U)},\bm{0}^{(U)},\bm{0}^{(U)}]$ for P2, where $u_{k}$, $d_{k}$ represent the index of the actions at $k^{th}$ decision node, $k=0,1,2,3$
@@ G.4 Joint-Perturbation Simultaneous Pseudo-Gradient (JPSPG)
The core idea in the JPSPG algorithm is the use of pseudo-gradient instead of computing the actual gradient of the utility to update players’ strategies. By perturbing the parameters of a utility function (which consists of the strategy), an unbiased estimator of the gradient of a smoothed version of the original utility function is obtained. Computing pseudo-gradient can often be cheaper than computing exact gradient, and at the same time suitable in scenarios where the utility (or objective) functions are "black-box" or unknown. Building on top of pseudo-gradient, martin2024joint proposed a method that estimates the pseudo-gradient with respect to all players’ strategies simultaneously. The implication of this is that instead of multiple calls to estimate the pseudo-gradient, we can estimate the pseudo-gradient in a single evaluation. More formally, let $\textbf{f}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ be a vector-valued function. Then, its smoothed version is defined as:
@@
In games with discrete-action spaces, where strategy is the probability distribution over the actions, JPSPG can be directly applied to get mixed strategy. However, for continuous-action games, a standard implementation would result in pure strategy solution than mixed. In order to compute a mixed strategy, we can turn into neural network as a strategy with an added randomness that can be learned as described in martin2023finding; martin2024joint. We similarly define two strategy networks for each player, the outputs of which are scaled based on the respective action bounds with the help of hyperbolic tangent (tanh) activation on the final layer. The input to the strategy networks (a single hidden layered neural network with 64 neurons and output neuron of action-space dimension) are the state of the player and a random variable whose mean and variance are trainable parameters. We follow the architecture as outlined by martin2024joint in their implementation of continuous-action Goofspiel. We would like to thank the authors for providing an example implementation of JPSPG on a normal-form game.
@@ G.6 Value Network Training Details
We briefly discuss the training procedure of the value networks. As mentioned in the main paper, both the primal and the dual value functions are convex with respect to $p$ and $\hat{p}$ respectively. As a result, we use Input Convex Neural Networks (ICNN) (amos2017icnn) as the neural network architecture. Starting from $T-\tau$, solutions of the optimization problem $\text{P}_{1}$ for sampled $(X,p)$ is saved and the convex value network is fit to the saved training data. The model parameters are saved and are then used in the optimization step at $T-2\tau$. This is repeated until the value function at $t=0$ is fit. The inputs to the primal value network are the joint states containing position and velocities of the players $X$ and the belief $p$.
@@
Appendix H Hyperparameter Sweep for PG Baselines
Here we report a sweep of hyperparameters across different learning rates and entropy coefficient for the PG MMD and PPO algorithms. Specifically, we run the algorithms with learning rates of $\{2.5e-5,2.5e-4,2.5e-2,2.5e-1\}$, and entropy coefficient of $\{0.01,0.05,0.1,0.2\}$. We also run RNaD with all four learning rates. The sweep is reported in Fig. 12.
Figure 12: Hyperparameter sweep for the baseline PG algorithms. RNaD, being a non-standard PG algorithm, doesn’t use entropy coefficient the way MMD and PPO do; hence, we copy the same plot across different entropy coefficient value for reference.
Appendix I Solving Hexner’s game via MPC
@@
Appendix J A Differentiable 11-vs-11 American Football Game
@@
J.1 Making Sense of Exploitability
While exploitability provides a good indication of whether or not the learned policy is converging to the equilibrium, without a well known baseline outcome, it is often difficult to make sense of how drastic the true nash equilibrium is compared to the resulting policy with non-zero $\epsilon$ exploitability. Hence, here we compare the resulting trajectories when each player switch to their best-response policy to get a sense of how “exploitable” are the strategies, qualitatively. We plot the comparison in Fig. 14. Note that unlike in Fig. 1(b), in Fig. 14 P2 plays a dual game, which results in slight change in P1’s policy.
Figure 14: Comparison of resulting behavior from the final P1 and P2’s policies vs their respective best-response (BR) policies. Both players’ behavior do not shift qualitatively when switching to their BR policies highlighting that the resulting policy with non-zero exploitability is indeed close to the equilibrium.